40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Message ID: 2611     Entry time: Thu Jan 21 15:25:13 2021
Author: aaron 
Type: DailyProgress 
Category: Lab Work 
Subject:  

Enter Thu Jan 21 15:25:08 2021

resealed the HEPA filter

over the lab's air intake with Loctite silicone sealant. The rubber cement I used last time delaminated, and I'm hoping this sealant contours to the rough wall paint and proves more effective.

Historically, sealing the HEPA filter decreased the lab temperature. Attachment 1 shows the lab temperature for the last 17 days -- as you can begin to see, sealing the HEPA actually increased the lab filter. This is unfortunate, but consistent with the HVAC blowing hot air despite being set to cool.

 

Organizing electronics rack

I relabelled the 'E' and 'W' path to 'North' and 'South'. This should be less confusing, as the lasers' fiber components run parallel from west-to-east (one laser is north of the other). It would be good to continue labeling, including the AC power cables. I started a techmart cart for cable solutions from digikey (+torque wrench, panel mount SMA feedthroughs, to avoid sending RF through BNC, rack mount components).

Bypassed lossy AOM

The South laser is losing >8dB passing through its AOM. I've taken this AOM out of the path, so the South laser now runs from its Faraday isolator directly to its EOM. For symmetry, I've also removed the North laser's EOM, so the north laser passes through an AOM before being sent to a 90-10 beam splitter. With this modification, the power at the laser launch for the north laser is 7.3 mW and the south is 6.2 mW (the AOM is expected to be lossier than the EOM).

For PSOMA we typically think of our signal being an amplitude modulation, while the pump is phase-modulated for locking. With this modification, the north laser is playing the role of 'signal', while the south laser plays the role of 'pump.' Fine for now, but of course eventually we'll want to swap in a higher power laser for the south (pump) path.

I noticed that the latex gloves being used to softly clamp the AOM have begun to stiffen and deteriorate. Touching the glove at all breaks it into small pieces. This is worrying, so I unmounted the AOM and replaced the gloves. I will look into butterfly mounts and soft, non-degradable cable clamps for the future.

miscellany

  • Pulled aside some bad BNC cables for repair (connector loose, no strain relief).

Exit Thu Jan 21 20:20:27 2021

Attachment 1: Screenshot_from_2021-01-21_19-57-47.png  38 kB  Uploaded Thu Jan 21 19:57:59 2021  | Hide | Hide all
Screenshot_from_2021-01-21_19-57-47.png
ELOG V3.1.3-